Angel’s Pizza delivers several types of pizza, which is sold for 499 pesos each and costs 300 pesos to make. The pizza shop has a policy that no charge will be asked if the delivery takes longer than a half an hour. Experience has shown that delivery takes longer than half an hour only 10% of the time. How much will the pizza shop expect to gain/lose in return?




1
Expert's answer
2022-03-07T14:39:01-0500


Let X be the random variable representing the charge that Angel’s Pizza gains or losses when it delivers a pizza. Given that the probability of making a unsuccessful delivery is 0.1, the probability of making a successful delivery will be equal to 1- 0.1 = 0.9


X P(X)

499 0.9

-300 0.1

E(X) = Σ XiPi = (499*0.9) + (-300*0.1) = 419.1

Answer: Since the expected value is positive, the pizza shop expects to gain 419.1 pesos in return